# Using thermodynamics to learn gravitational wave physics

### v1

@@ Abstract
Black holes are some of the most interesting objects in the universe. While they first arise in the complicated behavior of general relativity, the physical laws ruling their behavior are surprisingly simple. For example, one of the core facts about black holes is that their area never decreases, much alike the entropy in thermodynamics. In this note directed at introductory physics students and their instructors, we use this similarity to understand properties of black hole physics using standard techniques from an undergraduate course in thermal physics. We explore the never-decreasing nature of black hole area to obtain bounds on the energy emitted in a black hole merger (a calculation originally done by Hawking). We show how this allows us to think of black holes in manners very similar to heat engines, and how these ideas have been used in modern gravitational wave observatories to test general relativity. This allows a research-level topic to be discussed in introductory physics lectures.
@@ I Introduction
One of the most interesting predictions that has been recently tested through gravitational wave observations is the “area theorem” [11, 2]. This result by Hawking [8, 9] states that, under mild hypotheses, the total area of black holes in the universe never decreases. Hence, the black hole area behaves very similarly to the entropy of usual thermal systems, and we can learn aspects of black hole physics through a straightforward analogy with textbook thermodynamics. In fact, this analogy has now been recognized as a profound statement about the behavior of gravity and developed into the field of black hole thermodynamics. Studying it has taught us much about the inner workings of gravity, thermodynamics, and quantum theory. Our goal in this note is to explain how these ideas can emerge in black hole mergers, and how they have been used in modern gravitational wave observatories to test general relativity in the presence of strong gravitational fields.
@@ VI Discussion
As the technology of gravitational wave observatories improves—and as new observatories are constructed—the uncertainties in the date will become progressively smaller. This means that, with time, tests of the area theorem will allows us to learn whether the universe indeed respects Hawking’s prediction, or whether one of the assumptions behind the theorem is wrong. Therefore, studying whether the area theorem holds in black hole mergers is a way of better understanding how good general relativity is as a description of gravity, and what could lie beyond.
@@
This perspective, however, changed considerably about fifty years ago, when Hawking [10, *hawking1976ParticleCreationBlackErratum] showed that quantum effects can allow a black hole to have a nonvanishing temperature. This supported the view that the area of a black hole should literally be interpreted as its entropy, with the detailed correspondence being given by the Bekenstein–Hawking formula,
@@ Acknowledgements.
NAA and CCRE contributed equally to this work and share first authorship. Figure 1 was drawn using Asymptote [4, 7]. We thank Bruno Arderucio Costa for many pleasant and helpful conversations about this work. CCRE also thanks Artur Soares and João Macedo for pedagogical suggestions. The work of NAA was supported by the São Paulo Research Foundation (FAPESP) under grant 2025/05161-0. The work of CCRE was supported by the Conselho Nacional de Desenvolvimento Cíentifico (CNPq) under grant 131663/2025-9.

### v2

@@ Abstract
Black holes are some of the most interesting objects in the universe. While they first arise in the complicated behavior of general relativity, the physical laws ruling their behavior are surprisingly simple. For example, one of the core facts about black holes is that their area never decreases, much like the entropy in thermodynamics. In this note directed at introductory physics students and their instructors, we use this similarity to understand properties of black hole physics using standard techniques from an undergraduate course in thermal physics. We explore the never-decreasing nature of black hole area to obtain bounds on the energy emitted in a black hole merger (a calculation originally done by Hawking). We show how this allows us to think of black holes in manners very similar to heat engines, and how these ideas have been used in modern gravitational wave observatories to test general relativity. This allows a research-level topic to be discussed in introductory physics lectures.
@@ I Introduction
One of the most interesting predictions that has been recently tested through gravitational wave observations is the “area theorem” [9, 10, 2]. This result by Hawking [6, 7] states that, under mild hypotheses, the total area of black holes in the universe never decreases. Hence, the black hole area behaves very similarly to the entropy of usual thermal systems, and we can learn aspects of black hole physics through a straightforward analogy with textbook thermodynamics. In fact, this analogy has now been recognized as a profound statement about the behavior of gravity and developed into the field of black hole thermodynamics. Studying it has taught us much about the inner workings of gravity, thermodynamics, and quantum theory. Our goal in this note is to explain how these ideas can emerge in black hole mergers, and how they have been used in modern gravitational wave observatories to test general relativity in the presence of strong gravitational fields.
@@ VI Discussion
As the technology of gravitational wave observatories improves—and as new observatories are constructed—the uncertainties in the data will become progressively smaller. This means that, with time, tests of the area theorem will allows us to learn whether the universe indeed respects Hawking’s prediction, or whether one of the assumptions behind the theorem is wrong. Therefore, studying whether the area theorem holds in black hole mergers is a way of better understanding how good general relativity is as a description of gravity, and what could lie beyond.
@@
This perspective, however, changed considerably about fifty years ago, when Hawking [8, *hawking1976ParticleCreationBlackErratum] showed that quantum effects can allow a black hole to have a nonvanishing temperature. This supported the view that the area of a black hole should literally be interpreted as its entropy, with the detailed correspondence being given by the Bekenstein–Hawking formula,
@@ Acknowledgements.
NAA and CCRE contributed equally to this work and share first authorship. Figure 1 was drawn using Asymptote [bowman2008AsymptoteVectorGraphics, hammerlindl2004AsymptoteDescriptiveVector]. We thank Bruno Arderucio Costa for many pleasant and helpful conversations about this work. CCRE also thanks Artur Soares and João Macedo for pedagogical suggestions. The work of NAA was supported by the São Paulo Research Foundation (FAPESP) under grant 2025/05161-0. The work of CCRE was supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under grant 131663/2025-9.